Q1 T17. Two fuel gases are mixed together prior to being burned with air in a furnace to produce a flue gas with the following composition: Co2 7%, Co 1%, O2 7% & N2 85% (on a dry basis). The analyses of the fuel gases are as follows: Fuel Gas A: CH4 80%, N2 20% Fuel Gas B CH4 60%, C2H6 20%, N2 20% All compositions are expressed on a mole basis. Calculate a The mole ratio of Gas A to Gas B in the fuel gas mixture. b The percentage excess air used in the furnace
Answer #1

Ans:

a) Given:

Fuel A- 80% CH4 , 20% N2

Fuel B- 60 % CH4, 20% C2H6 , 20% N2

Step 1: Convert all the masses into moles.

Let the total mass of both fuel gases be 100 g each.

So in fuel gas A: let the mass of CH4 =80 g

And the mass of N2 =20 g

Now, no. Of moles= given mass / molar mass

So no. of moles of CH4 in fuel gas = 80 g / 16g = 5 moles

And no. Of moles of N2 in fuel gas A= 20g/ 28g = 0.714 moles

Total no. Of moles of fuel gas A= 5.714

In fuel gas B:

Let the mass of CH4= 60g,

Mass of C2H6= 20g

& mass of N2= 20 g

Now, no. Of moles of CH4 in fuel gas B= 60g/16g = 3.75 mol

No. Of moles of C2H6 in fuel gas B= 20g / 30g = 0.666 moles

& no. Of moles of N2= 20g/ 28g = 0.714 moles

Total no. Of moles of fuel gas B= 5.13 moles

Now, mole ratio of gas A to gas B= 5.714/ 5.13 =1.11

B) Excess air: The amount of excess air within the system can be determined by analyzing the amount of O2 in the fuel gas.

Thus the CO level in the fuel gas is assumed to be very low and incomplete combustion is neglected, hence the formula used to calculate excess air on a dry basis is:

Excess air %=( 0.895× A%) / (0.21 - A%)

A% is the percentage of Oxygen in fuel gas mixture.

A% = 7% =0.07

Therefore, Excess air %= 0.895 × 0.07 /(0.21-0.07)

= 0.4475

= 44.75%